OpenStudy (princesskitty01):

Will anyone show me how to do this please What is the simplified form of

alones (alones):

Form of?

OpenStudy (iam.paige):

Of What ?

OpenStudy (princesskitty01):

hang on

OpenStudy (princesskitty01):

http://prntscr.com/cq6xs0

alones (alones):

So you need to simplify \[\frac{ 3 }{ x^2\ }~\over ~\frac{ 1 }{ x^3 }\]

OpenStudy (princesskitty01):

yes

OpenStudy (princesskitty01):

x over 3 1 over 3x 3x 3 over x

alones (alones):

Hrmm alright so we can't really simplify \[\frac{ 3 }{ x^2}\] neither we cannot simpliy the second equation. So line them up \[\frac{ 3 }{ x^2}\ and \frac{ 1 }{ x^3}\] \(\ 3*1=3\) and there's no exponents left. So the real answer would be just \(\ 3x\).

OpenStudy (princesskitty01):

okay thx

alones (alones):

Yeah you welcome (: